# Adversarial Attacks on Data Attribution

### v2

@@ Abstract
Data attribution aims to quantify the contribution of individual training data points to the outputs of an AI model, which has been used to measure the value of training data and compensate data providers. Given the impact on financial decisions and compensation mechanisms, a critical question arises concerning the adversarial robustness of data attribution methods. However, there has been little to no systematic research addressing this issue. In this work, we aim to bridge this gap by detailing a threat model with clear assumptions about the adversary’s goal and capabilities and proposing principled adversarial attack methods on data attribution. We present two methods, Shadow Attack and Outlier Attack, which generate manipulated datasets to inflate the compensation adversarially. The Shadow Attack leverages knowledge about the data distribution in the AI applications, and derives adversarial perturbations through “shadow training”, a technique commonly used in membership inference attacks. In contrast, the Outlier Attack does not assume any knowledge about the data distribution and relies solely on black-box queries to the target model’s predictions. It exploits an inductive bias present in many data attribution methods—outlier data points are more likely to be influential—and employs adversarial examples to generate manipulated datasets. Empirically, in image classification and text generation tasks, the Shadow Attack can inflate the data-attribution-based compensation by at least $200\%$, while the Outlier Attack achieves compensation inflation ranging from $185\%$ to as much as $643\%$.
@@ 3.2 The adversary
The two proposed attack methods in Section 4 and Section 5 respectively rely on 2 and 3, but do not depend on both simultaneously.
@@ 4.2 Adversarial manipulation by maximizing shadow compensation rate
Given the shadow compensation rate, a natural idea is to apply adversarial perturbations to $Z$ by solving $Z_{1}^{a}=\operatorname*{arg\,max}\nolimits_{Z^{\prime}\in\mathcal{N}(Z)}c_{s}(Z^{\prime})$, where $\mathcal{N}(Z)$ specifies the space of datasets with undetectable perturbations. In practice, however, solving this optimization problem has two technical challenges. Firstly, the objective $c_{s}(\cdot)$ is discrete and can be difficult to optimize. Secondly, when doing iterative-style optimization, such as gradient ascent, $\tau^{(i)}$’s need to be updated and re-evaluated at every step as it depends on $Z$ and hence $M^{(i)}=T(Z^{(i)}\cup Z)$. This leads to two problems: On the one hand, updating $M^{(i)}$’s requires retraining, which is computationally heavy; on the other hand, many data attribution methods are computationally expensive, even with all the data and (retrained) models available. Hence, maximizing $c_{s}(\cdot)$ requires repeatedly retraining and running data attributions on perturbed datasets, which may be infeasible for many cases.
@@ 6.1 Experimental setup
We consider two machine-learning tasks: image classification and text generation. For image classification, we experiment on MNIST (LeCun, 1998), Digits (Jiang et al., 2023) and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, with different target models including Logistic Regression (LR), Multi-layer Perceptron (MLP), Convolutional Neural Networks (CNN), and ResNet-18 (He et al., 2016). We also employ three popular data attribution methods, including Influence Function (Koh & Liang, 2017), TRAK (Park et al., 2023), and Data Shapley (Ghorbani & Zou, 2019). For text generation, we conduct experiments on NanoGPT (Karpathy, 2022) trained on the Shakespeare dataset (Karpathy, 2015), with TRAK as the data attribution method. Finally, for the image classification settings, we evaluate both Shadow Attack and Outlier Attack, while for the text generation setting, we evaluate Outlier Attack only as 2 usually does not hold for generative AI settings. The experimental settings are summarized in Table 1.
@@ 6.4 Baseline reference
6.5 Visualization of the perturbation
Finally, Figure 3 provides visualizations of two examples of the adversarial perturbed images CIFAR-10. In both cases, the perturbations are barely visible to human eyes, however, they are highly effective in terms of the success of the attack.
(a) Original.
Influential for 00 validation data points.
(b) Shadow Attack.
Influential for $75$ validation data points.
(c) Outlier Attack.
Influential for $105$ validation data points.
(d) Original.
Influential for $1$ validation data points.
(e) Shadow Attack.
Influential for $38$ validation data points.
(f) Outlier Attack.
Influential for $29$ validation data points.

### v3

@@ Abstract
Data attribution aims to quantify the contribution of individual training data points to the outputs of an AI model, which has been used to measure the value of training data and compensate data providers. Given the impact on financial decisions and compensation mechanisms, a critical question arises concerning the adversarial robustness of data attribution methods. However, there has been little to no systematic research addressing this issue. In this work, we aim to bridge this gap by detailing a threat model with clear assumptions about the adversary’s goal and capabilities and proposing principled adversarial attack methods on data attribution. We present two methods, Shadow Attack and Outlier Attack, which generate manipulated datasets to inflate the compensation adversarially. The Shadow Attack leverages knowledge about the data distribution in the AI applications, and derives adversarial perturbations through “shadow training”, a technique commonly used in membership inference attacks. In contrast, the Outlier Attack does not assume any knowledge about the data distribution and relies solely on black-box queries to the target model’s predictions. It exploits an inductive bias present in many data attribution methods—outlier data points are more likely to be influential—and employs adversarial examples to generate manipulated datasets. Empirically, in image classification and text generation tasks, the Shadow Attack can inflate the data-attribution-based compensation by at least $200\%$, while the Outlier Attack achieves compensation inflation ranging from $185\%$ to as much as $643\%$. Our implementation is ready at https://github.com/TRAIS-Lab/adversarial-attack-data-attribution.
@@ 3.2 The adversary
The two proposed attack methods in Section 4 and Section 5 respectively rely on 2 and 3, but do not depend on both simultaneously. As a result, the first method is a gray-box attack method while the second method is a black-box attack method.
@@ 4.2 Adversarial manipulation by maximizing shadow compensation rate
Given the shadow compensation rate, the natural idea is to apply adversarial perturbations to $Z$ by solving $Z_{1}^{a}=\operatorname*{arg\,max}\nolimits_{Z^{\prime}\in\mathcal{N}(Z)}c_{s}(Z^{\prime})$, where $\mathcal{N}(Z)$ specifies the space of datasets with undetectable perturbations. In practice, however, solving this optimization problem has two technical challenges. Firstly, the objective $c_{s}(\cdot)$ is discrete and can be difficult to optimize. Secondly, when doing iterative-style optimization, such as gradient ascent, $\tau^{(i)}$’s need to be updated and re-evaluated at every step as it depends on $Z$ and hence $M^{(i)}=T(Z^{(i)}\cup Z)$. This leads to two problems: On the one hand, updating $M^{(i)}$’s requires retraining, which is computationally heavy; on the other hand, many data attribution methods are computationally expensive, even with all the data and (retrained) models available. Hence, maximizing $c_{s}(\cdot)$ requires repeatedly retraining and running data attributions on perturbed datasets, which may be infeasible for many cases.
@@ 6 Experiments
In this section, we evaluate the effectiveness of the proposed attack methods by the increase of compensation share after manipulating the data with the proposed attack methods.
@@ 6.1 Experimental setup
We consider two machine-learning tasks: image classification and text generation. For image classification, we experiment on MNIST (LeCun, 1998), Digits (Jiang et al., 2023) and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, with different target models including Logistic Regression (LR), Multi-layer Perceptron (MLP), Convolutional Neural Networks (CNN), and ResNet-18 (He et al., 2016). We also employ three popular data attribution methods, including Influence Function (Koh & Liang, 2017), TRAK (Park et al., 2023), and Data Shapley (Ghorbani & Zou, 2019). For text generation, we conduct experiments on NanoGPT (Karpathy, 2022) trained on the Shakespeare dataset (Karpathy, 2015), with TRAK as the data attribution method. Finally, for the image classification settings, we evaluate both Shadow Attack and Outlier Attack, while for the text generation setting, we evaluate Outlier Attack only as 2 usually does not hold for generative AI settings. For the data attribution algorithms, we adopt the implementation from the dattri library (Deng et al., 2024). The experimental settings are summarized in Table 1.
@@ B.3 Attack methods
Appendix C Additional experiments
In this section, we conduct additional experiments under various settings, including ablation studies and more.
C.1 Ablation study: changing value of $|Z_{0}|$ and $|Z_{0}!|$
We test the performance of our methods under different sizes of $|Z_{0}|$ and $|Z_{1}|$ to understand the effect of data set sizes. For settings (a), (c), and (d) in Table 1, we consider decreasing $(|Z_{0}|,|Z_{1}|)$ from $(10000,11000)$ to $(5000,6000)$ and increasing $(|Z_{0}|,|Z_{1}|)$ from $(10000,11000)$ to $(15000,16000)$, experimenting with both the Shadow Attack and the Outlier Attack. For the setting (e), i.e., the text generation setup, the original $|Z_{0}|$ was already small (originally, $(|Z_{0}|,|Z_{1}|)=(4706,6274)$), which takes only $30\%$ of the data. Hence, we consider increasing the data size to two different extents. Note that since the Shadow Attack is infeasible in this setting, only the Outlier Attack’s results are shown for setting (2). The results for the two attack methods are respectively shown in Tables 5 and 6. We see that the results demonstrate the effectiveness of our methods across different sizes of the dataset. Overall, with a few exceptions, the Ratio of the Compensation Share appears to become even higher when the dataset size is larger.
Table 5: Results of Shadow Attack for various $(|Z_{0}|,|Z_{1}|)$ settings.
Setting $(|Z_{0}|,|Z_{1}|)$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
(a) $(5000,6000)$ $0.0153$ $0.0764$ $499.3\%$ $0.996$ $0.003$ $0.001$
$(10000,11000)$ $0.0098$ $0.0477$ $456.1\%$ $0.955$ $0.038$ $0.007$
$(15000,16000)$ $0.0050$ $0.0373$ $746.0\%$ $0.963$ $0.036$ $0.001$
(c) $(5000,6000)$ $0.0168$ $0.0539$ $320.7\%$ $0.857$ $0.129$ $0.014$
$(10000,11000)$ $0.0112$ $0.0467$ $417.0\%$ $0.781$ $0.195$ $0.024$
$(15000,16000)$ $0.0002$ $0.0062$ $3100.0\%$ $0.431$ $0.568$ $0.001$
(d) $(5000,6000)$ $0.0206$ $0.0413$ $200.7\%$ $0.696$ $0.174$ $0.130$
$(10000,11000)$ $0.0095$ $0.0213$ $217.3\%$ $0.655$ $0.259$ $0.086$
$(15000,16000)$ $0.0057$ $0.0092$ $161.5\%$ $0.264$ $0.616$ $0.120$
Table 6: Results of Outlier Attack for various $(|Z_{0}|,|Z_{1}|)$ settings.
Setting $(|Z_{0}|,|Z_{1}|)$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
(a) $(5000,6000)$ $0.0153$ $0.0747$ $488.9\%$ $0.998$ $0.002$ $0.000$
$(10000,11000)$ $0.0098$ $0.0631$ $643.9\%$ $0.980$ $0.017$ $0.003$
$(15000,16000)$ $0.0050$ $0.0400$ $800.0\%$ $0.964$ $0.036$ $0.000$
(c) $(5000,6000)$ $0.0168$ $0.0334$ $198.8\%$ $0.720$ $0.241$ $0.039$
$(10000,11000)$ $0.0112$ $0.0668$ $596.4\%$ $0.799$ $0.173$ $0.028$
$(15000,16000)$ $0.0002$ $0.0051$ $2550.0\%$ $0.397$ $0.603$ $0.000$
(d) $(5000,6000)$ $0.0206$ $0.0411$ $199.5\%$ $0.761$ $0.133$ $0.106$
$(10000,11000)$ $0.0095$ $0.0176$ $185.2\%$ $0.562$ $0.354$ $0.084$
$(15000,16000)$ $0.0057$ $0.0219$ $384.2\%$ $0.731$ $0.192$ $0.077$
(e) $(4706,6274)$ $0.0031$ $0.0035$ $262.9\%$ $0.392$ $0.461$ $0.147$
$(7843,9411)$ $0.0016$ $0.0064$ $400.0\%$ $0.420$ $0.507$ $0.073$
$(12549,14116)$ $0.0029$ $0.0214$ $737.9\%$ $0.400$ $0.543$ $0.057$
C.2 Ablation study: data attribution method
Next, we conduct an ablation study by varying the data attribution methods used in the evaluation under the experiment setting (b) in Table 1. Specifically, apart from Data Shapley, we consider two additional data attribution methods, Influence Function, and TRAK, for evaluating the compensation share, experimenting with both the Shadow Attack and the Outlier Attack. Note that the Adversary has no knowledge about what data attribution method will be used by the AI Developer. The results are shown in Tables 7 and 8. The results below show that the proposed attack methods are highly effective for all three data attribution methods, as reflected by the Ratio of the Compensation Share.
Table 7: Results of Shadow Attack with different data attribution methods of setting (b).
Attribution Method $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
Data Shapley $0.0352$ $0.0152$ $0.0435$ $286.2\%$ $0.533$ $0.333$ $0.134$
Influence Function $0.0352$ $0.0496$ $0.1004$ $202.4\%$ $0.980$ $0.000$ $0.020$
TRAK $0.0352$ $0.0392$ $0.0936$ $238.8\%$ $0.820$ $0.100$ $0.080$
Table 8: Results of Outlier Attack with different data attribution methods of setting (b).
Attribution Method $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
Data Shapley $0.0250$ $0.0112$ $0.0218$ $194.6\%$ $0.440$ $0.380$ $0.180$
Influence Function $0.0250$ $0.0186$ $0.0442$ $237.6\%$ $0.920$ $0.060$ $0.020$
TRAK $0.0250$ $0.0160$ $0.0412$ $257.5\%$ $0.780$ $0.180$ $0.040$
C.3 White-box attack
In this section, we further consider white-box attacks as an oracle reference, where the Adversary has full knowledge of the target model and has access to the model parameters. For the choice of attack method under the white-box threat model, we experiment with the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015) for setting (a) and the Projected Gradient Descent(PGD) (Mkadry et al., 2017) for setting (d). The results are shown in Table 9. Overall, compared to this oracle white-box attack, the proposed black-box attacks are only slightly worse in terms of the Ratio of the Compensation Share. This further confirms that the proposed methods are highly effective.
Table 9: Results of white-box attacks under settings (a) and (d).
Setting Attack Method Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
(a) FGSM $0.0098$ $0.0649$ $662.2\%$ $0.990$ $0.007$ $0.003$
(d) PGD $0.0095$ $0.0222$ $233.7\%$ $0.689$ $0.204$ $0.107$
C.4 Adversarial attack under data augmentation
In this section, we test the performance of our proposed attack methods when the AI Developer utilize data augmentation techniques when training the model. Specifically, after the AI Developer gathers data, the training dataset is then formed by a combination of the gathered data and its augmented version. For simplicity, we consider settings (a), (c), and (d) in Table 1 and consider standard image augmentation methods such as random cropping and random affine transformation to images. With data augmentation, the compensation share will be attributed back to the original data point if its augmented version is identified as influential. The results are shown in Tables 10 and 11. Compared to the original setting without data augmentation, we cannot draw a definite conclusion on whether data augmentation helps defend the proposed attacks since the trend is unclear. However, overall, we can conclude that the proposed attacks are still highly effective even when data augmentation is utilized by the AI Developer.
Table 10: Results of Shadow Attack under data augmentation
Setting $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
(a) $0.0098$ $0.0099$ $0.0209$ $211.1\%$ $0.632$ $0.244$ $0.124$
(c) $0.0098$ $0.0084$ $0.0516$ $614.3\%$ $0.894$ $0.095$ $0.011$
(d) $0.0098$ $0.0091$ $0.0747$ $820.9\%$ $0.959$ $0.035$ $0.006$
Table 11: Results of Outlier Attack under data augmentation.
Setting $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
(a) $0.0098$ $0.0099$ $0.0287$ $289.9\%$ $0.776$ $0.158$ $0.066$
(c) $0.0098$ $0.0084$ $0.0691$ $822.6\%$ $0.853$ $0.113$ $0.034$
(d) $0.0098$ $0.0091$ $0.0481$ $528.6\%$ $0.916$ $0.065$ $0.019$
C.5 Compensating only correct prediction
In this section we consider the setup when the AI Developer only rewards training data points that are highly influential to validation samples with correct predictions. In detail, after training the model and before calculating of compensation share, the AI Developer will test the model on a private validation set, select those validation points that are correctly predicted, and only attribute compensation to influential training points for these validation points. We test this on settings (a), (c), and (d). Note that the Fraction of Change metric is no longer feasible in this setup since the model trained on the original and manipulated training set may not share the same validation set. The results are shown in Table 12. Overall, our proposed attacks are still highly effective based on the Ratio of Compensation Share.
Table 12: Results when compensating only correctly predicted data points.
Attack Type Setting $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share
Original Manipulated Ratio
Shadow Attack (a) $0.0098$ $0.0050$ $0.0382$ $764.0\%$
(c) $0.0098$ $0.0090$ $0.0752$ $835.6\%$
(d) $0.0098$ $0.0110$ $0.0353$ $320.9\%$
Outlier Attack (a) $0.0098$ $0.0050$ $0.0273$ $546.0\%$
(c) $0.0098$ $0.0090$ $0.0552$ $613.3\%$
(d) $0.0098$ $0.0110$ $0.0227$ $206.4\%$
C.6 Adversarial attack on large-scale setup
In this section, we scale up our experiment and see how the proposed attacks generalize in this large-scale experiment. Specifically, we consider the ResNet-18 model (He et al., 2016) on the Tiny ImageNet dataset (Le & Yang, 2015) with $|Z_{0}|,|Z_{1}|)=(50000,60000)$, and we perturb $|Z_{1}^{a}|=100$ training points. Both the Shadow Attack and Outlier Attack are tested, and the results are shown in Table 13. It is evident that from the Ratio of Compensation Share, two proposed attacks are still highly effective in a large-scale experiment.
Table 13: Results of Shadow Attack and Outlier Attack under large-scale setup .
Attack Type $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
Outlier Attack 0.0017 $0.0009$ $0.0056$ $622.2\%$ $0.456$ $0.498$ $0.046$
Shadow Attack 0.0017 $0.0009$ $0.0181$ $2011.1\%$ $0.855$ $0.131$ $0.014$
Appendix D Visualization of the perturbation
Finally, Figure 3 provides visualizations of two examples of the adversarial perturbed images from MNIST and CIFAR-10. In both cases, the perturbations are barely visible to human eyes, however, they are highly effective in terms of the success of the attack.
(a) Original.
Influential for 00 validation data points.
(b) Shadow Attack.
Influential for $75$ validation data points.
(c) Outlier Attack.
Influential for $105$ validation data points.
(d) Original.
Influential for $1$ validation data points.
(e) Shadow Attack.
Influential for $38$ validation data points.
(f) Outlier Attack.
Influential for $29$ validation data points.